Tierra, Nico, and Alex had $865 altogether. Tierra spent 2/5 of her money. Nico spent $40 , and Alex spent twice as much as Tierra. If the three friends had the same amount of money left, how much money did Alex have in the beginning.
Answer #1
This exact question was answered earlier today.

See link.

http://www.jiskha.com/display.cgi?id=1298550296
answered by: Shanelle
Answer #3
If Tierra spent 2/5 and Alex spent twice it means Alex spends 4/5. So now we subtract the excess to work with the fractions. So 865-40= $825. So now add the fractions to see the parts you have so 2/5+ 4/5=6/5. So we find 6/5 of 825=990 then we find the total so 990+40=$1030 so to find for the 3 parts u divide. 1030÷3=$343.33
answered by: Aidan
Answer #4
Tierra in total will have 5 parts Alex would have 7 parts And nico will have 3 parts with the remainder of 40 So 865-40=$825 5+7+3=15 parts total 825÷15=55 for 1 part 55×7=$385 Alex =$385 This is correct and it is teacher proven This is the answer for it, answer sheet has shown it and teachers have also confirmed it
answered by: Hayley Jittan